Frans88:

If 0.329 g of Fe3O4 and 0.812 g of Al are available for the following reaction: [4] 3 Fe3O4 + 8 Al → 4 Al2O3 + 9 Fe a) What would be the limiting reactant?

Vocaloid:

The equation is already balanced (always check first) so you can move to the stoichiometry. To calculate the limiting reactant, determine whether 0.329g Fe3O4 *or* 0.812g Al makes less product (I'll choose Fe as the product) I'll show you the calculation for Fe3O4 and let you do the Al calculation on your own. \[\frac{ 0.329g~Fe_{3}O_{4} }{ }\frac{ 1~mol~Fe_{3}O_{4} }{ 231.55g }\frac{ 9~mol~Fe }{ 3~mol~Fe_{3}O_{4} }\] notice how we go from g Fe3O4, divide by the molar mass of Fe3O4 to get moles Fe3O4. then, using the ratio of the equation, we divide by 3 moles Fe3O4 and multiply by 9 moles Fe to get moles Fe. Repeat this calculation with 0.812g Al. Whichever reactant makes **less** Fe is the limiting reactant.

Frans88:

Thank you very much, I got it now.